# Hessian Aware Low-Rank Weight Perturbation for Continual Learning

### v2

@@ 5.2 Empirical Accuracy
We provide the average accuracies on the six benchmarks in Table II and Table III. From these numerical results, we can conclude our method achieved state-of-the-art performances compared to the baseline methods. From these numerical results, we can conclude our method achieved state-of-the-art performances compared to the baseline methods.
According to the evaluations on CIFAR-100 Split and CIFAR-100 Superclass in Table II, we can observe that our proposed HALRP outperforms the recent methods (e.g., GPM (replay-based), APD and WSN (expansion-based)) with a significant margin (e.g., with an improvement about $1\%\sim 3\%$). Especially, we also evaluated the performances of all the methods under different amounts of training data (i.e., $5\%\sim 100\%$) on these two benchmarks. It was interesting to see that our proposed method had advantages to deal with extreme cases with limited data. Compared to other methods, the performance margins become more significant with fewer training data. For example, on CIFAR-100 Split, our proposed HALRP outperforms APD and WSN with an improvement of $1.28\%$ and $1.36\%$ respectively, but these margins will dramatically rise to $\sim 8\%$ if we reduce the training set to $5\%$ of the total data. These results showed that our method can work better under these extreme cases.
On Five-dataset, we adopted two kinds of neural networks, i.e., AlexNet and ResNet18, to verify the effectiveness of HALRP under different backbones. We can observe that our method consistently outperforms the other methods under different backbones, indicating the applicability and flexibility for different model choices. Compared with the analyses in the following part, we can conclude that our method is also robust under different orders of the tasks.
@@
On TinyImageNet dataset, we evaluated our methods on two settings, one with 20-split and another with 40-split, and the resuls are posted in Table III(d). According to the emprirical results, we can conclude that our proposed method can perform well on this challenging dataset, with a better trade-off between the average accuracy and task order robustness. Compared to the previous methods APD and APDfix that aim to address the issue of task order robustness, our HALRP can perform well with a more consistent manner. Furthermore, our methods also have the advantages regarding the computational efficiency and memory consumption that we will discuss later in the following sections.
Additionally, we also provide comparisons with IRU [33] that is also based on low-rank decomposition. We note that the official code of IRU only contains a demo for multi-layer perceptron (MLP), and no complete code for the convolutional network (even LeNet) can be found. Due to this limitaton, we cannot implement IRU under our setting. To make fair comparisons, we re-implement our method under the setting of IRU [33]. We describe the setting as follows: 1) Dataset Protocol: [33] generated 20 random tasks on PMNIST, rather than 10 tasks in this work. Moreover, [33] randomly divided the CIFAR100 dataset into 20 tasks, rather than the superclass-based splitting in our paper. Thus, we denote the two datasets in [33] as “PMNIST-20” and “CIFAR100-20”, to distinguish from the “PMNIST” and “CIFAR100-Superclass” used in the common setting. (2) Model Architecture: Apart from the architectures we introduced before, we also follow the MLP used in [33], a three-layer (fully-connected) multilayer perceptron with 256 hidden nodes. The results of our method and IRU are listed in Table IV. We can conclude that our method achieved a large performance gain compared to IRU.
@@ 5.3 Robustness on Task Orders
We evaluate the robustness of these algorithms under different task orders. Following the protocol of [9], we assessed the task order robustness with the Order-normalized Performance Disparity (OPD) metric, which is computed as the disparity between the performance $\bar{P}_{t}^{R}$ of task $t$ on $R$ random task sequences: $\mathrm{OPD}_{t}\triangleq\max\{\bar{P}_{t}^{1},\dots,\bar{P}_{t}^{R}\}-\min\{\bar{P}_{t}^{1},\dots,\bar{P}_{t}^{R}\}$. The maximum OPD (MOPD) and average OPD (AOPD) are defined by
$\mathrm{MOPD}$ $\triangleq\max\{\mathrm{OPD}_{1},\dots,\mathrm{OPD}_{T}\}$ (14)
$\mathrm{AOPD}$ $\triangleq\frac{1}{T}\sum_{t=1}^{T}\mathrm{OPD}_{t}$
@@
Thus, smaller AOPD and MOPD indicate better task-order robustness. For CIFAR100-Split and Superclass datasets, we adopted different orders by following [9]. For the P-MNIST and Five-dataset, we report the averaged results over five different task orders. The results on MOPD and AOPD on P-MNIST, Five-dataset, Omniglot-Rotation, TinyImageNet, and CIFAR-100 Splits/Superclass are reported in Table III(a), Table III(b), Table III(c), Table III(d) and Table V, respectively. According to these results, we can generally conclude that our method can achieve a better accuracy-robustness trade-off, compared to the recent baseline APD.
CIFAR-100 Split (with LeNet) CIFAR-100 SuperClass (with LeNet)
@@
TABLE V: Task order robustness evaluation on CIFAR-100 Split/Superclass with different amounts of training data.
@@ 5.4 Handling the Catastrophic Forgetting
We then evaluated the abilities of different methods for overcoming catastrophic forgetting, a long-standing issue in continual learning. We illustrate the average forgetting (BWT) on the CIFAR100 Split/SuperClass dataset under different amounts of data in Fig. 2. Our proposed method can effectively address the forgetting issue, comparable with some state-of-the-art methods like WSN. Apart from the averaged results shown in Fig. 2, we further provide detailed forgetting statistics within five different task orders in Appendix E.2.
@@ 5.5 Model Increment Analysis
Another potential issue in continual learning is the model size growth as the number of tasks increases. To compare increased model capacities among different methods, we visualize the ratio of increased parameters w.r.t to base model on the sequential 20 tasks of CIFAR100-Superclass dataset in Fig. 3(a). We can observe that our proposed HALRP can better control the model capacity increment. After the first four tasks, the model parameters are increased by around $19\%$ and grow slowly during the last sixteen tasks (finally $\leq 28\%$). In contrast, the model parameters of APD grew quickly by around $40\%$ in the first two or three tasks and keep a high ratio during the following tasks.
@@ 5.6 Computational Efficiency
We visualize the average time complexity ratio on the ten tasks of the PMNIST dataset in Fig. 3(b). The time complexity ratio is computed by dividing the accumulated time across the tasks w.r.t. the time cost on the first task of single-task learning. We observe that our proposed method is also computational efficient limited time consumption compared to most baselines, with a better trade-off between performance and efficiency. Especially, our method needs less training time compared to WSN, BMKP and APD.
The major reason that should account for the inefficiency of WSN is that this method needs to optimize a weight scores for each model parameter and then generate binary masks by locating a certain sparsity quantile among the weight scores in each layer, which is very time-consuming. Especially, we can observe a non-linear increase behaviour for BMKP and APD in Fig. 3(b). In BMKP, the pattern basis of the core knowledge space increased along the sequential tasks and then the knowldege projection between two memory levels will cost more time as the tasks pass. A sever drawback of APD method is that when the task $t$ arrives for learning, it needs to recover the parameters $\theta^{\star}_{i}$ for all previous tasks $i=0,1,...,t-1$ and then apply the regularization $\Sigma_{i=0}^{t-1}\|\theta_{t}-\theta^{\star}_{i}\|^{2}_{2}$ between the weights of current task (i.e., $\theta_{t}$) and each previous task (i.e., $\theta^{\star}_{i}$). This defect becomes more time-consuming as more tasks arrive. Besides, an extra $k$-means clustering process is needed for the hierarchical knowledge consolidation among tasks in APD. The difference of the computational efficiency become more significant on the challenging scenarios with more tasks and complex network architectures.
@@ 5.7 Ablation Studies
Ablation study about the rank selection We report the additional studies to evaluate the impacts of the hyperparameters selection on the rank selection process. We conduct the following ablations: (1) varying the warm-up epochs $n_{r}$, (2) varying the loss approximation rate $\alpha$, (3) LRP: omitting the importance estimation (step 6 of Algorithm 1), and (4) Random LRP: replacing step 6 of Algorithm 1 with a random decomposition. The relevant results on CIFAR100-Splits are depicted in Table VI. We can observe that the performance will not significantly rise as $n_{r}$ increases, which indicates that only one epoch is enough for warm-up training. When decreasing the approximate rate $\alpha$, the accuracy will decay. Furthermore, if we apply the random decomposition of the model, the accuracy will sharply drop, indicating the necessity of our Hessian aware-decomposition procedure.
Effect of the regularization coefficients $\lambda_{0}$ and $\lambda_{1}$ To investigate the effects of the regularization coefficients introduced in Eq. 12, we further conducted experiments on two datasets: CIFAR100-Superclass and CIFAR100-Splits. We adopted values from $\{0,1e-6,1e-5,1e-4,1e-3\}$ for $\lambda_{0}$ and $\{5e-4,1e-4,5e-5\}$ for $\lambda_{1}$. For each combination, we reported MOPD/AOPD for task order robustness as well as the average accuracy on the above two datasets, and the results are illustrated in Fig. 4. We can see that the regularization coefficients $\lambda_{0}$ and $\lambda_{1}$ have potential effect on the final accuracy and task order robustness. In our experiments, we choose the optimal hyperparameters by validation. Furthermore, some selected hyperparameters are also applied to other baselines to realize fair comparisons, e.g., we set $L_{2}$-regularizer coefficient $\lambda_{1}=1e-4$ for all other methods on CIFAR100-Superclass and CIFAR100-Splits. See Appendix F.3 for more details.
@@ Appendix B Proof to Theorem 1 and Discussion
Assume that a neural network of $L$ layers with vectorized weights $(\boldsymbol{\omega}^{\star}_{1},\dots,\boldsymbol{\omega}^{\star}_{L})$ that have converged to local optima, such that the first and second order optimality conditions are satisfied, i.e., the gradient is zero, and the Hessian is positive semi-definite. Suppose a perturbation $\Delta\boldsymbol{\omega}^{\star}_{1}$ applied to the first layer weights, then we have the loss increment
$\begin{split}\mathcal{L}(\boldsymbol{\omega}^{\star}_{1}-\Delta\boldsymbol{\omega}^{\star}_{1},&\dots,\boldsymbol{\omega}^{\star}_{L})-\mathcal{L}(\boldsymbol{\omega}^{\star}_{1},\dots,\boldsymbol{\omega}^{\star}_{L})\\ &\leq\|\frac{1}{2}\mathbf{H}_{1}\|_{F}\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F}+o(\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F}),\end{split}$ (15)
where $\mathbf{H}_{1}=\Delta\mathcal{L}(\boldsymbol{\omega}^{\star}_{1})$ is the Hessian matrix at only the variables of the first layer weights.
@@
$\begin{split}\mathcal{L}(\boldsymbol{\omega}^{\star}_{1}&-\Delta\boldsymbol{\omega}^{\star}_{1},\dots,\boldsymbol{\omega}^{\star}_{L})-\mathcal{L}(\boldsymbol{\omega}^{\star}_{1},\dots,\boldsymbol{\omega}^{\star}_{L})\\ &=\frac{1}{2}\Delta\boldsymbol{\omega}^{\star T}_{1}\mathbf{H}_{1}\Delta\boldsymbol{\omega}^{\star}_{1}+o(\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F})\\ &\leq\|\frac{1}{2}\Delta\boldsymbol{\omega}^{\star T}_{1}\mathbf{H}_{1}\Delta\boldsymbol{\omega}^{\star}_{1}+o(\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F})\|_{F}\\ &\leq\|\frac{1}{2}\Delta\boldsymbol{\omega}^{\star T}_{1}\mathbf{H}_{1}\Delta\boldsymbol{\omega}^{\star}_{1}\|_{F}+o(\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F})\\ &\leq\|\frac{1}{2}\mathbf{H}_{1}\|_{F}\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F}+o(\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F}).\end{split}$
@@ E.2 Performance on Alleviate Forgetting on Different Task Orders
In the main paper, we report the forgetting performance on the average of five task orders (A-E) on CIFAR-100 Splits and Superclass. In this part, we provide the performance on forgetting under each task order in Fig. 5 and Fig. 6, under different amounts (i.e., $5\%\sim 100\%$) of training data from.
@@ F.1 Dataset Preparation
CIFAR-100 Splits/SuperClass We used the CIFAR-100 Splits/Superclass dataset following the evaluation protocol of [9], and follow the task order definition of [9] to test the algorithms with five different task orders (A-E).
For the CIFAR-100 Split, the task orders are defined as:
@@
For the CIFAR-100 Superclass, the task orders are:
@@
Dataset CIFAR-100 Split CIFAR-100 Super PMNIST Five dataset Omniglot TinyImageNet 40-split TinyImageNet 20-split

### v5

@@ 5.2 Empirical Accuracy
We provide the average accuracies on the six benchmarks in Table II and Table III. From these numerical results, we can conclude our method achieved state-of-the-art performance compared to the baseline methods.
According to the evaluations on CIFAR100-Split and CIFAR100-SuperClass in Table II, we can observe that our proposed HALRP outperforms the recent methods (e.g., GPM (replay-based), APD and WSN (expansion-based)) with a significant margin (e.g., with an improvement about $1\%\sim 3\%$). Especially, we also evaluated the performances of all the methods under different amounts of training data (i.e., $5\%\sim 100\%$) on these two benchmarks. It was interesting to see that our proposed method had advantages in dealing with extreme cases with limited data. Compared to other methods, the performance margins become more significant with fewer training data. For example, on CIFAR100-Split, our proposed HALRP outperforms APD and WSN with an improvement of $1.28\%$ and $1.36\%$ respectively, but these margins will dramatically rise to $\sim 8\%$ if we reduce the training set to $5\%$ of the total data. These results showed that our method can work better in these extreme cases.
On Five-dataset, we adopted two types of neural network, i.e., AlexNet and ResNet18, to verify the effectiveness of HALRP under different backbones. We can observe that our method consistently outperforms the other methods under different backbones, indicating the applicability and flexibility for different model choices. Compared with the analyses in the following part, we can conclude that our method is also robust under different orders of tasks.
@@
On TinyImageNet dataset, we evaluated our methods on two settings, one with 20-split and another with 40-split, and the results are posted in Table III(d). According to the empirical results, we can conclude that our proposed method can perform well on this challenging dataset, with a better trade-off between the average accuracy and task order robustness. Compared to the previous methods APD and APDfix that aim to address the issue of task order robustness, our HALRP can perform well in a more consistent manner. Furthermore, our methods also have advantages regarding computational efficiency and memory consumption that we will discuss later in the following sections.
Additionally, we also provide comparisons with IRU [34] that is also based on low-rank decomposition. We note that the official code of IRU only contains a demo for multi-layer perceptron (MLP), and no complete code for the convolutional network (even LeNet) can be found. Due to this limitation, we cannot implement IRU under our setting. To make fair comparisons, we re-implement our method under the setting of IRU [34]. We describe the setting as follows: 1) Dataset Protocol: [34] generated 20 random tasks on PMNIST, rather than 10 tasks in this work. Moreover, [34] randomly divided the CIFAR100 dataset into 20 tasks, rather than the superclass-based splitting in our paper. Thus, we denote the two datasets in [34] as “PMNIST-20” and “CIFAR100-20”, to distinguish them from the “PMNIST” and “CIFAR100-SuperClass” used in the common setting. (2) Model Architecture: Apart from the architectures we introduced before, we also follow the MLP used in [34], a three-layer (fully-connected) multilayer perceptron with 256 hidden nodes. The results of our method and IRU are listed in Table IV. We can conclude that our method achieved a large performance gain compared to IRU.
@@ 5.3 Robustness on Task Orders
We evaluate the robustness of these algorithms under different task orders. Following the protocol of [9], we assessed the task order robustness with the Order-normalized Performance Disparity (OPD) metric, which is computed as the disparity between the performance $\bar{P}_{t}$ of task $t$ under $R$ different task orders: $\mathrm{OPD}_{t}\triangleq\max\{\bar{P}_{t}^{1},\dots,\bar{P}_{t}^{R}\}-\min\{\bar{P}_{t}^{1},\dots,\bar{P}_{t}^{R}\}$. The maximum OPD (MOPD) and average OPD (AOPD) are defined by
$\mathrm{MOPD}$ $\triangleq\max\{\mathrm{OPD}_{0},\dots,\mathrm{OPD}_{T-1}\}$ (14)
$\mathrm{AOPD}$ $\triangleq\frac{1}{T}\sum_{t=0}^{T-1}\mathrm{OPD}_{t}$
@@
Thus, smaller AOPD and MOPD indicate better task-order robustness. For CIFAR100-Split and SuperClass datasets, we adopted different orders by following [9]. For the P-MNIST and Five-dataset, we report the averaged results over five different task orders. The results on MOPD and AOPD on P-MNIST, Five-dataset, Omniglot-Rotation, TinyImageNet, and CIFAR100-Splits/-SuperClass are reported in Table III(a), Table III(b), Table III(c), Table III(d) and Table V, respectively. According to these results, we can generally conclude that our method can achieve a better accuracy-robustness trade-off, compared to the recent baseline APD.
CIFAR100-Split (with LeNet) CIFAR100-SuperClass (with LeNet)
@@
TABLE V: Task order robustness evaluation on CIFAR100-Split/SuperClass with different amounts of training data.
@@ 5.4 Handling the Catastrophic Forgetting
We then evaluated the abilities of different methods for overcoming catastrophic forgetting. We illustrate the average forgetting (BWT) on the CIFAR100 Split/SuperClass dataset under different amounts of data in Fig. 2. Our proposed method can effectively address the forgetting issue, comparable with some state-of-the-art methods like WSN. We further provide detailed forgetting statistics within five different task orders in Appendix E.2.
@@ 5.5 Model Increment Analysis
Another potential issue in continual learning is the model size growth as the number of tasks increases. To compare increased model capacities among different methods, we visualize the ratio of increased parameters w.r.t to base model on the sequential 20 tasks of CIFAR100-SuperClass dataset in Fig. 3(a). We can observe that our proposed HALRP can better control the model capacity increment. After the first four tasks, the model parameters increased by around $19\%$ and grow slowly during the last sixteen tasks (finally $\leq 28\%$). In contrast, the model parameters of APD grew quickly by around $40\%$ in the first two or three tasks and kept a high ratio during the following tasks.
@@ 5.6 Computational Efficiency
We visualize the average time complexity ratio on the ten tasks of the PMNIST dataset in Fig. 3(b). The time complexity ratio is computed by dividing the accumulated time across the tasks w.r.t. the time cost on the first task of single-task learning. We observe that our proposed method is also computationally efficient with limited time consumption compared to most baselines, with a better trade-off between performance and efficiency. Especially, our method needs less training time compared to WSN, BMKP and APD.
The major reason that should account for the inefficiency of WSN is that this method needs to optimize a weight score for each model parameter and then generate binary masks by locating a certain sparsity quantile among the weight scores in each layer, which is very time-consuming. Especially, we can observe a non-linear increase in behavior for BMKP and APD in Fig. 3(b). In BMKP, the pattern basis of the core knowledge space increases along the sequential tasks and then the knowledge projection between two memory levels will cost more time as the tasks pass. A severe drawback of APD method is that when the task $t$ arrives for learning, it needs to recover the parameters $\theta^{\star}_{i}$ for all previous tasks $i=0,1,...,t-1$ and then apply the regularization $\Sigma_{i=0}^{t-1}\|\theta_{t}-\theta^{\star}_{i}\|^{2}_{2}$ between the weights of current task (i.e., $\theta_{t}$) and each previous task (i.e., $\theta^{\star}_{i}$). This defect becomes more time-consuming as more tasks arrive. Besides, an extra $k$-means clustering process is needed for the hierarchical knowledge consolidation among tasks in APD. The difference of computational efficiency becomes more significant on the challenging scenarios with more tasks and complex network architectures. As for PRD, it is inefficient as this method converges slower than other methods, due to the reason that it adopted supervised contrastive loss instead of cross-entropy loss.
@@ 5.7 Ablation Studies
Ablation study about the rank selection We report the additional studies to evaluate the impacts of hyperparameter selection on the rank selection process. We conduct the following ablations: (1) varying the warm-up epochs $n_{r}$, (2) varying the loss approximation rate $\alpha$, (3) LRP: omitting the importance estimation (step 6 of Algorithm 1), and (4) Random LRP: replacing step 6 of Algorithm 1 with a random decomposition. The relevant results on CIFAR100-Splits are depicted in Table VI. We can observe that the performance will not significantly rise as $n_{r}$ increases, which indicates that only one epoch is enough for warm-up training. When decreasing the approximate rate $\alpha$, the accuracy will decay. Furthermore, if we apply the random decomposition of the model, the accuracy will sharply drop, indicating the necessity of our Hessian aware-decomposition procedure.
Effect of the regularization coefficients $\lambda_{0}$ and $\lambda_{1}$ To investigate the effects of the regularization coefficients introduced in Eq. 12, we further conducted experiments on two datasets: CIFAR100-SuperClass and CIFAR100-Splits. We adopted values from $\{0,1e-6,1e-5,1e-4,1e-3\}$ for $\lambda_{0}$ and $\{5e-4,1e-4,5e-5\}$ for $\lambda_{1}$. For each combination, we reported MOPD/AOPD for task order robustness as well as the average accuracy on the above two datasets, and the results are illustrated in Fig. 4. We can see that the regularization coefficients $\lambda_{0}$ and $\lambda_{1}$ have potential effects on the final accuracy and task order robustness. In our experiments, we choose the optimal hyperparameters by validation. Furthermore, some selected hyperparameters are also applied to other baselines to realize fair comparisons, e.g., we set $L_{2}$-regularizer coefficient $\lambda_{1}=1e-4$ for all other methods on CIFAR100-SuperClass and CIFAR100-Splits. See Appendix F.3 for more details.
@@ Appendix B Proof to Theorem 1 and Discussion
Assume that a neural network of $L$ layers with vectorized weights $(\boldsymbol{\omega}^{\star}_{1},\dots,\boldsymbol{\omega}^{\star}_{L})$ that have converged to local optima, such that the first and second order optimality conditions are satisfied, i.e., the gradient is zero, and the Hessian is positive semi-definite. Suppose a perturbation $\Delta\boldsymbol{\omega}^{\star}_{1}$ applied to the first layer weights, then we have the loss change
$\begin{split}|\mathcal{L}(\boldsymbol{\omega}^{\star}_{1}-\Delta\boldsymbol{\omega}^{\star}_{1},&\dots,\boldsymbol{\omega}^{\star}_{L})-\mathcal{L}(\boldsymbol{\omega}^{\star}_{1},\dots,\boldsymbol{\omega}^{\star}_{L})|\\ &\leq\frac{1}{2}\|\mathbf{H}_{1}\|_{F}\cdot\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F}+o(\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F}),\end{split}$ (15)
where $\mathbf{H}_{1}=\nabla^{2}\mathcal{L}(\boldsymbol{\omega}^{\star}_{1})$ is the Hessian matrix at only the variables of the first layer weights.
@@
$\begin{split}|\mathcal{L}(\boldsymbol{\omega}^{\star}_{1}&-\Delta\boldsymbol{\omega}^{\star}_{1},\dots,\boldsymbol{\omega}^{\star}_{L})-\mathcal{L}(\boldsymbol{\omega}^{\star}_{1},\dots,\boldsymbol{\omega}^{\star}_{L})|\\ &=|\frac{1}{2}\Delta\boldsymbol{\omega}^{\star T}_{1}\mathbf{H}_{1}\Delta\boldsymbol{\omega}^{\star}_{1}+o(\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F})|\\ &\leq\|\frac{1}{2}\Delta\boldsymbol{\omega}^{\star T}_{1}\mathbf{H}_{1}\Delta\boldsymbol{\omega}^{\star}_{1}\|_{F}+o(\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F})\\ &\leq\frac{1}{2}\|\mathbf{H}_{1}\|_{F}\cdot\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F}+o(\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F}).\end{split}$
@@ E.2 Performance on Alleviate Forgetting on Different Task Orders
In the main paper, we report the forgetting performance on the average of five task orders (A-E) on CIFAR100 Splits and SuperClass. In this part, we provide the performance on forgetting under each task order in Fig. 5 and Fig. 6, under different amounts (i.e., $5\%\sim 100\%$) of training data from.
@@ F.1 Dataset Preparation
CIFAR100 Splits/SuperClass We used the CIFAR100 Splits/SuperClass dataset following the evaluation protocol of [9], and follow the task order definition of [9] to test the algorithms with five different task orders (A-E).
For the CIFAR100 Split, the task orders are defined as:
@@
For the CIFAR100 SuperClass, the task orders are:
@@
Dataset CIFAR100 Split CIFAR100 Super PMNIST Five dataset Omniglot TinyImageNet 40-split TinyImageNet 20-split